please help me with this word problem. I just cant figure out how to set it up.
sally's age plus 8 is equal to gart's age doubled. the sum of their ages is 124. how old is sally?

Answer by richwmiller(17219)

s+8=2g
s=2g-8
s+g=124
s=124-g
2g-8=124-g
3g=132
g=44
s+8=2g
s+8=88
s=80
check
44+88=132
ok